# What We Talk About When We Talk About LMs: Implicit Paradigm Shifts and the Ship of Language Models

### v2

@@ Abstract
The term Language Models (LMs), as a time-specific collection of models of interest, is constantly reinvented, with its referents updated much like the Ship of Theseus replaces its parts but remains the same ship in essence. In this paper, we investigate this Ship of Language Models problem, wherein scientific evolution takes the form of continuous, implicit retrofits of key existing terms. We seek to initiate a novel perspective of scientific progress, in addition to the more well-studied emergence of new terms. To this end, we construct the data infrastructure based on recent NLP publications. Then, we perform a series of text-based analyses toward a detailed, quantitative understanding of the use of Language Models as a term of art. Our work highlights how systems and theories influence each other in scientific discourse, and we call for attention to the transformation of this Ship that we all are contributing to.
@@ 1 Introduction
In this paper, we seek a quantitative description of a field’s continuous evolution. More specifically, we inquire into the Ship of LMs paradox, i.e., the aforementioned reconstruction of the term Language Models. We decipher this evergreen term’s rapidly changing referents, contexts, and usages across time. We develop a semi-automatic, generalizable framework to extract and organize two closely related sets of keywords: (1) mentions of the collective LM concept, and (2) specific model names, and construct a dataset of 7,650 papers from the 10 most recent, major NLP conferences.
@@
Our work highlights the astonishing extent of change subtly encoded in the seemingly unchanged overarching terms. We hope the Ship of LMs can serve as a new perspective to understand the field’s progress, and that our methodology can serve as an entry point for additional finer-grained measurements of subtle changes in rapidly growing fields.
@@ 2 Related Work
Various studies review the history of NLP conferences and the ACL Anthology Hall et al. (2008); Anderson et al. (2012); Bollmann et al. (2023), as well as the community that contributed to the field’s trajectory Abdalla et al. (2023); Movva et al. (2023). Other works identified the field’s transition points and themes. Hou et al. (2019) proposes an automatic framework to extract key entities (tasks, dataset, etc.). Uban et al. (2021) explores a similar goal via topic modeling, and Pramanick et al. (2023) further identified such entities that causally shaped the field’s important stages. More recently, there has also been a specific focus on the changes brought by LLMs Min et al. (2023); Fan et al. (2023); Zhao et al. (2023) and the impact on the related communities Saphra et al. (2023); Liang et al. (2024). Aside from text-based analysis, interviews and surveys (Gururaja et al., 2023; Michael et al., 2023, etc.) have also provided valuable qualitative insights for the disruptive shifts.
@@ 3.2.1 Notations
As described, we seek to build two related collections of key entities, one marking the mentions of LMs as a general term, and the other marking specific model names. The two are respectively denoted by $\mathcal{L}$ (from LMs) and $\mathcal{M}$ (from Models).
@@
since “language model” is the substring of most of its subcategories, e.g., “large language models”, “Korean language models”, or “language modeling”, and searching for “language model” [^4] Here we include other upper-case forms, e.g., “Language Models”. This is implemented similar to Aliases in §3.2.2. covers all such variations. We also include the most common acronyms, “LLM” (Large Language Models) and “PLM” (Pre-trained Language Models). The construction of $\mathcal{M}$ is elaborated in §3.2.2.
@@ 3.2.2 Constructing $\mathcal{M}$ from the text
To compose the final list, we sorted names following a simple heuristic (illustrated in Appendix A.2). When we encounter a new model $m$ not in $\mathcal{M}$, we add $m$ to $\mathcal{M}$ and initialize its alias list as $[m]$. Additional aliases of an entry $m_{j}$ are appended to its list. Variations of existing entries are recorded but not added as an alias, and candidates that are not the name of a model (e.g., BLEU) are discarded.
For each entry of $m$, we also manually retrieved the original paper or documentation to determine if there is an explicit dependency on another model. In all, $\mathcal{M}$ has a total of 97 model entries and 146 aliases. With the two keyword lists, $\mathcal{L}$ and $\mathcal{M}$, we are ready to examine how LMs are resolved and extract diachronic patterns.
Our dataset and code will be open-sourced, and we provide the full list of models involved and the LLM prompts in Appendix A.3 and A.4.
@@ 4.1 Wind in the Sails: Surging Mentions, Speeding Conclusions
We consider the average $N^{\mathcal{L}}$ of all papers at a certain conference, written as $\bar{N^{\mathcal{L}}}$. With the division of LM-related and non-related papers, $\bar{N^{\mathcal{L}}}$ can be calculated as the (weighted) average of the two groups. If the above assumption holds true, $\bar{N^{\mathcal{L}}}$ can be estimated as follows: On the one hand, non-LM-related papers always have $N^{\mathcal{L}}=0$ by definition. On the other hand, within the LM-related group, the average of $N^{\mathcal{L}}$ would also be stable across conferences since the use of the term LMs remains generally unchanged as hypothesized. As both groups – constituting the full set of papers at a conference when combined – see no qualitative change in their within-group $N^{\mathcal{L}}$, the determining factor for the overall average $\bar{N^{\mathcal{L}}}$ is the ratio of the two groups’ size (as the weight), i.e., the percentage of LM-related papers. Thus, we can draw an estimate from the scale of the first data point ($\bar{N^{\mathcal{L}}}=4.29$ for $35\%$ LM-related at ACL 2020). For instance, 54% of papers at EMNLP 2021 are LM-related, which is 1.54×1.54$\times$1.54 × that of ACL 2020 (35%). We can thus scale $\bar{N^{\mathcal{L}}}$ with the same ratio, $1.54\times 4.29\approx 6.56$, as the estimate for EMNLP 2021.
@@
However, we see a strong deviation from the estimated growth starting 2022. $\bar{N^{\mathcal{L}}}$ has since been on a super-exponential growth, eventually being 80% higher than estimated at ACL 2023 and 168% higher at EMNLP 2023 (where $\bar{N^{\mathcal{L}}}$ nearly doubled in just half a year). In other words, the Ship is not just sailing better (more papers), but it is also undergoing reconstruction (referents of LM are changing). The distinct patterns pre- and post-2022 despite a similar background increase highlight the necessity to study the Ship of LMs as a dynamic concept, as emergence of a term is not sufficient for mining the deeper nuances as such.
@@
Figure 3: Compositions of model names in the main conference papers at EMNLP 2020 (left) and 2023 (right), arranged counter-clock-wise by their size. Branches indicate dependency, e.g., GPT-3.5 is based on and fine-tuned from GPT-3. Each group of models with the same root is represented by a unique color shared between graphs.
@@
First, we see evidence supporting our prior observations on the patterns of $N^{\mathcal{L}}$. Earlier conferences form a cluster where no significant difference is noted; yet, starting 2022, every conference has a significantly higher ${N^{\mathcal{L}}}$ than most or even all of its predecessors. However, the distributions of $N$ tell a distinct story. For most pairs, there is little or no evidence for a difference in distribution. There also isn’t a similar line that divides the earlier and most recent conferences. For instance, the distribution of $N$ for EMNLP in 2023 does not have a significant difference with that in 2020, despite all its specialties. In fact, we even see an opposite case: conferences in 2022 and 2023 – the exact time of the super-linear boosts of $\bar{N^{\mathcal{L}}}$ – contain significantly fewer model mentions than before. In other words, we arrive at the conclusions faster: the information conveyed via specific models has not increased, but more is drawn about LMs collectively. Various factors might account for this trend: It might be that we are increasingly proficient in maintaining the Ship, or that more weight is put on the holistic concept of the Ship compared with closer examinations on the level of specific models, etc. More effort would be needed to discern the exact cause, and deeper understandings of the local context of LMs (e.g. similar to Jurgens et al., 2018) would be especially relevant for future work.
@@ 4.2 Oak, Pine, or Cedar Planks: Which models are we talking about?
How much remains as the replacement of earlier components goes on? We calculate the Jaccard similarity between compositions of conferences to quantify how much is shared across any two conferences, shown in Figure 4. We observe that Jaccard similarity between conferences monotonically decreases for subsequent conferences, which matches the Ship of LMs case where its parts change over time. For two consecutive conferences, the Jaccard similarity is usually only 71% to 85%; the index quickly drops to 45% to 57% with an interval of just two years, and to 24% to 31% in three years. We also note that dissimilarity is rapidly increasing, with EMNLP 2023 sharing a 52% Jaccard similarity with ACL 2023 just half a year ago, 42% with EMNLP in 2022, and $\leq$38with all other predecessors. With the representing models thoroughly reshuffled in as short as $<$5 years, the “shelf life” of our conclusions and knowledge of LM has seen a new low, thus bringing unprecedented challenges for long-term tasks and literature studies.
@@
Figure 5 displays the proportion of publications with absolute majority components for the full data (left) and the top 25% with the highest $N^{\mathcal{L}}$ (right). The most notable components are marked with the same color as in Fig. 3. Other models are collected as the grey bar (“others”), and the proportion with no absolute majority is denoted with a striped pattern. Around 80% of the papers contain an absolute majority model. Specifically, we get a glance at the astounding traction of BERT before more recent paradigm shifts: it dominated up to 60.5% of all papers and 68.6% of the most LM-centered ones.
Interestingly, more focus on the collective LM terms did not entail a more balanced composition. In fact, they are often more biased: The percentage of papers with an absolute majority is higher in the most LM-centered quartile than overall for all 7 conferences before EMNLP 2022. There has also been a fresh wind, however. In the most recent 3 conferences, we see fewer dominant components when a paper focuses heavily on the generalized LM terms. More importantly, the chance of an absolute majority in both groups has been on a continuing decline and both reached an unprecedentedly lower level: 70.5% for all papers and 67% for the top quarter. We call to keep monitoring the heterogeneity (or lack thereof) in LM papers given the (still) high presence of major components and a visibly surging presence of the GPT component at the most recent EMNLP 2023.
@@ 4.3 Lembos or Trireme: Factoring in context
The extent to which a paper focuses on LMs implies different use scenarios. For instance, a work may utilize and mention them for data processing but doesn’t concern the science of LMs per se. This usually implies lower $N^{\mathcal{L}}$ in contrast to another work that studies an emergent property of LLMs. As we develop our understanding of how LMs are embodied in papers, we need to consider how the composition of papers factors into their usage.
We rank the LM-related papers at a conference by $N^{\mathcal{L}}$ and compare the ones with the most and least focus on the LM terms. Specifically, we extract the top and bottom quarters, denoted as ${Q4}^{+}$ and ${Q1}^{-}$ respectively. Figure 6 compares compositions of $N$ from ${Q4}^{+}$ and ${Q1}^{-}$ of EMNLP 2023 and shows the differences in the 10 largest components. In ${Q4}^{+}$, the GPT component covers an additional 13.2% of $N$ compared with ${Q4}^{+}$, followed by LLaMA (6.5%) and T5 (5.0%). Meanwhile, the BERT component takes as much as an extra 23.6% in ${Q1}^{-}$, where the roles of the LMs terms are most peripheral.
@@
Figure 7: Differences in the model composition of EMNLP 2020 and 2023, measured respectively on the most LM-focused quarter (${Q4}^{+}$; darker colors) and the least one (${Q1}^{-}$; lighter colors). Blue/Red denotes an increase/decrease of share. We annotate beside each bar the exact numbers in 2020 and 2023, separated by ‘/’.
We further measure how the same group changes over time, comparing model compositions at an early conference (EMNLP 2020) versus a most recent one (EMNLP 2023). We depict the change of 15 largest components in Figure 7. The findings are consistent: the most LM-centered group adapts thoroughly to the latest models, while ${Q1}^{-}$ sees a much more modest change, or even increased interest for models pre-2020 like BART Lewis et al. (2020) and COMET Bosselut et al. (2019). Most notable is the contrast for BERT: For ${Q4}^{+}$, the once-dominant component loses as much as 37% of $N$, going from 55.8% to 19%. However, in ${Q1}^{-}$, the BERT component remains almost untouched; it still firmly takes up about 42% of $N$ despite all the new models. This demonstrates that the trend in less LM-centered groups is not merely a moderated or delayed version of that in the top ones, but indeed represents a distinct interpretation and resolution of LMs.
The findings converge to an interesting division: the newest components or those with the latest major models (GPT-3, ChatGPT, LlaMA, and Flan-T5, inter alia; all of which are post-2022) instantly enter the most LM-centered discourse, while the least LM-centered ones persist to favor certain earlier models for a longer period. Thus, we should be aware of the chasm between the co-occurring yet distinct contexts that eventually map to different constructions of the same Ship. For instance, a novel property found with 2023 models as default could be problematic if communicated to a BERT-centered subfield without regard, and further used to justify the use of LMs in a new stake despite wildly different ship drafts and capabilities. In return, the context difference may further hinder communications between groups representing the varied use and interpretations of LMs.

### v3

@@ Abstract
The term Language Models (LMs) as a time-specific collection of models of interest is constantly reinvented, with its referents updated much like the Ship of Theseus replaces its parts but remains the same ship in essence. In this paper, we investigate this Ship of Language Models problem, wherein scientific evolution takes the form of continuous, implicit retrofits of key existing terms. We seek to initiate a novel perspective of scientific progress, in addition to the more well-studied emergence of new terms. To this end, we construct the data infrastructure based on recent NLP publications. Then, we perform a series of text-based analyses toward a detailed, quantitative understanding of the use of Language Models as a term of art. Our work highlights how systems and theories influence each other in scientific discourse, and we call for attention to the transformation of this Ship to which we all are contributing. [^1] The data and code of this work is available at https://github.com/CurlyZhu/Ship_of_LMs.
@@ 1 Introduction
In this paper, we seek a quantitative description of a field’s continuous evolution. More specifically, we inquire into the Ship of LMs paradox, i.e., the aforementioned reconstruction of the term Language Models. We decipher this evergreen term’s rapidly changing referents, contexts, and usages across time. We develop a semi-automatic, generalizable framework to extract and organize two closely related sets of keywords: (1) mentions of the collective LM concept, and (2) specific model names, and construct a dataset of 7,650 papers from the 10 most recent major NLP conferences.
@@
Our work highlights the astonishing extent of change subtly encoded in the seemingly unchanged overarching terms. We hope the Ship of LMs can serve as a new perspective to understand the field’s progress, and that our methodology can serve as an entry point for finer-grained measurements of subtle changes in rapidly growing scientific fields.
@@ 2 Related Work
Various studies review the history of NLP conferences and the ACL Anthology Hall et al. (2008); Anderson et al. (2012); Bollmann et al. (2023), as well as the community that contributed to the field’s trajectory Abdalla et al. (2023); Movva et al. (2024). Other works identified the field’s transition points and themes: Hou et al. (2019) proposes an automatic framework to extract key entities (tasks, dataset, etc.); Uban et al. (2021) explores a similar goal via topic modeling; and Pramanick et al. (2023) further identified such entities that causally shaped the field’s important stages. More recently, there has also been a specific focus on the changes brought by LLMs Min et al. (2023); Fan et al. (2024); Zhao et al. (2023) and the impact on the related communities Saphra et al. (2024); Liang et al. (2024). Aside from text-based analysis, interviews and surveys (Gururaja et al., 2023; Michael et al., 2023, etc.) have also provided valuable qualitative insights for the disruptive shifts.
@@ 3.1 Dataset Construction
Figure 1: The full pipeline for constructing the model dictionaries (§3.2.2). The LLM agent follows a formatted prompt (§A.4) to automatically identify potential model names. The extracted strings are merged and ranked by frequency to form the list of candidate names. Then, the authors manually validate whether it is a new entry, an alias, or other based on a fixed protocol: (1) whether the candidate is indeed a valid model name, (2) whether it refers to the same model as an existing entry, and (3) whether it’s already covered by an alias of that entry.
@@ 3.2.1 Notations
As described, we seek to build two related collections of key entities, one marking the mentions of LMs as a general term, and the other marking specific model names. The two are respectively denoted by $\mathcal{L}$ (from LM) and $\mathcal{M}$ (from Models).
@@
since “language model” is the substring of most of its subcategories, e.g., “large language models”, “Korean language models”, or “language modeling”, and thus searching for “language model” [^5] Here we include other upper-case forms, e.g., “Language Models”. This is implemented similar to Aliases in §3.2.2. covers all such variations. We also include the most common acronyms, “LLM” (Large Language Models) and “PLM” (Pre-trained Language Models). The construction of $\mathcal{M}$ is elaborated in §3.2.2.
@@ 3.2.2 Constructing $\mathcal{M}$ from the text
To compose the final list, we manually validate and register the name candidates following a simple heuristic (Figure 1). When we encounter a new model $m$ not in $\mathcal{M}$, we add $m$ to $\mathcal{M}$ and initialize its alias list as $[m]$. Additional aliases of an entry $m_{j}$ are appended to its list. Variations of existing entries are recorded but not added as an alias, and candidates that are not the name of a model (e.g., BLEU) are discarded.
For each entry of $m$, we also manually retrieved the original paper or documentation to determine if there is an explicit dependency on another model. In all, $\mathcal{M}$ has a total of 103 model entries and 155 aliases. With the two keyword lists, $\mathcal{L}$ and $\mathcal{M}$, we are ready to examine how LMs are resolved and extract diachronic patterns.
Our dataset and code is available at https://github.com/CurlyZhu/Ship_of_LMs. We also provide more details of the implementation in the Appendix: the full list of models involved (A.2), the setup of the model name detector (A.3), and the LLM prompts (A.4).
@@ 4 Experiments and Findings
Year (prefix) 20- 21- 22- 23-
Venue (suffix) A E N A E A N E A E
Total 778 750 477 709 846 700 442 827 1074 1047
LM-related 273 337 239 333 454 407 282 571 794 877
Table 1: Total numbers of conference papers in this work after pre-processing. We denote the conferences henceforth with their first letters (ACL, EMNLP, and NAACL), and denote the years as prefixes for brevity. For instance, ‘21-E’ represents EMNLP 2021.
@@ 4.1 Wind in the Sails: Surging Mentions, Speeding Conclusions
We consider the average $N^{\mathcal{L}}$ of all papers at a certain conference, written as $\bar{N^{\mathcal{L}}}$. With the division of LM-related and non-related papers, $\bar{N^{\mathcal{L}}}$ can be calculated as the (weighted) average of the two groups. If the above assumption holds true, $\bar{N^{\mathcal{L}}}$ can be estimated as follows: On the one hand, non-LM-related papers always have $N^{\mathcal{L}}=0$ by definition. On the other hand, within the LM-related group, the average of $N^{\mathcal{L}}$ would also be stable across conferences since the use of the term LMs remains generally unchanged as hypothesized. As both groups – constituting the full set of papers at a conference when combined – see no qualitative change in their within-group $N^{\mathcal{L}}$, the determining factor for the overall average $\bar{N^{\mathcal{L}}}$ is the ratio of the two groups’ size (as the weights in weighted average), i.e., the percentage of LM-related papers. Thus, we can draw an estimate from the scale of the first data point ($\bar{N^{\mathcal{L}}}=4.29$ for $35\%$ LM-related at ACL 2020). For instance, 54% of papers at EMNLP 2021 are LM-related, which is 1.54×1.54$\times$1.54 × that of ACL 2020 (35%). We can thus scale $\bar{N^{\mathcal{L}}}$ with the same ratio, $1.54\times 4.29\approx 6.56$, as the estimate for EMNLP 2021.
@@
However, we see a strong deviation from the estimated growth starting 2022. $\bar{N^{\mathcal{L}}}$ has since been on an exponential growth, eventually being 80% higher than estimated at ACL 2023 and 168% higher at EMNLP 2023 (where $\bar{N^{\mathcal{L}}}$ nearly doubled in just half a year). In other words, the Ship is not just sailing better (more papers), but it is also undergoing reconstruction (referents of LM are changing). The distinct patterns pre- and post-2022 despite a similar background increase highlight the necessity to study the Ship of LMs as a dynamic concept, as emergence of a term is not sufficient for mining the deeper nuances as such.
@@
Figure 4: Compositions of model names in the main conference papers at EMNLP 2020 (left) and 2023 (right), arranged counterclockwise by their size. Branches indicate dependency, e.g., GPT-3.5 is based on and fine-tuned from GPT-3. Each group of models with the same root is represented by a unique color shared between graphs.
@@
First, we see evidence supporting our prior observations on the patterns of $N^{\mathcal{L}}$. Earlier conferences form a cluster where no significant difference is noted; yet, starting 2022, every conference has a significantly higher ${N^{\mathcal{L}}}$ than most or even all of its predecessors. However, the distributions of $N$ tell a distinct story. For most pairs, there is little or no evidence for a difference in distribution. There also isn’t a similar line that divides the earlier and most recent conferences. For instance, the distribution of $N$ for EMNLP in 2023 does not have a significant difference with that in 2020, despite all its specialties. In fact, we even see an opposite case: conferences in 2022 and 2023 – the exact time of the super-linear boosts of $\bar{N^{\mathcal{L}}}$ – contain significantly fewer model mentions than before. In other words, we arrive at the conclusions “faster”: the information conveyed via specific models has not increased, but more is drawn about LMs collectively. Various factors might account for this trend: It might be that we are increasingly proficient in maintaining the Ship, or that more weight is put on the holistic concept of the Ship compared with closer examinations on the level of specific models, etc. More effort would be needed to discern the exact cause, and a deeper understanding of the local context of LMs (e.g., similar to the work of Jurgens et al., 2018) would be especially relevant for future work.
@@ 4.2 Oak, Pine, or Cedar Planks: Which models are we talking about?
How much remains as the replacement of earlier components goes on? We calculate the Jaccard similarity between compositions of conferences to quantify how much is shared across any two conferences, shown in Figure 5. We observe that Jaccard similarity between conferences monotonically decreases for subsequent conferences, which matches the Ship of LMs case where its parts change over time. For two consecutive conferences, the Jaccard similarity is usually only 71% to 86%; the index quickly drops to 45% to 56% with an interval of just two years, and to 24% to 31% in three years. We also note that the dissimilarity is rapidly increasing, with EMNLP 2023 sharing a 52% Jaccard similarity with ACL 2023 just half a year ago, 42% with EMNLP in 2022, and $\leq$38with all other predecessors. With the representing models thoroughly reshuffled in as short as $<$5 years, the “shelf life” of our conclusions and knowledge of LM has seen a new low, thus bringing unprecedented challenges for long-term tasks and literature studies.
@@
Figure 6 displays the proportion of publications with absolute majority components for the full data (left) and the top 25% with the highest $N^{\mathcal{L}}$ (right). The most notable components are marked with the same color as in Fig. 4. Other models are collected as the grey bar (“others”), and the proportion with no absolute majority is denoted with a striped pattern. Around 80% of the papers contain an absolute majority model. Specifically, we get a glance at the astounding traction of BERT before more recent paradigm shifts: it dominated up to 61.1% of all papers and 67.9% of the most LM-centered ones.
Interestingly, more focus on the collective LM terms did not entail a more balanced composition. In fact, they are often more biased: The percentage of papers with an absolute majority is higher in the most LM-centered quartile than overall for all 7 conferences before EMNLP 2022. There has also been a fresh wind, however. In the most recent 3 conferences, we see fewer dominant components when a paper focuses heavily on the generalized LM terms. More importantly, the chance of an absolute majority in both groups has been on a continuing decline and both reached an unprecedentedly lower level: 70.2% for all papers and 66.3% for the top quarter. We call to keep monitoring the heterogeneity (or lack thereof) in LM papers given the (still) high presence of major components and a visibly surging presence of the GPT component at the most recent EMNLP 2023.
@@ 4.3 Lembos or Trireme: Factoring in context
The extent to which a paper focuses on LMs implies different use scenarios. For instance, a work may utilize and mention them for data processing but doesn’t concern the science of LMs per se. This usually implies lower $N^{\mathcal{L}}$ in contrast to, say, another work that studies an emergent property of LLMs. As we explore how LMs are embodied in papers, we need to consider how the composition of papers factors into the usage of “LMs”.
We rank the LM-related papers at a conference by $N^{\mathcal{L}}$ and compare between the ones with the most and least focus on the LM terms. Specifically, we extract the top and bottom quarters, denoted as ${Q4}^{+}$ and ${Q1}^{-}$ respectively. Figure 7 compares compositions of $N$ from ${Q4}^{+}$ and ${Q1}^{-}$ of EMNLP 2023 and shows the differences in the 10 largest components. In ${Q4}^{+}$, the GPT component covers an additional 12.7% of $N$ compared with ${Q1}^{-}$, followed by LLaMA (6.5%) and T5 (5.2%). Meanwhile, the BERT component takes as much as an extra 24.0% in ${Q1}^{-}$, where the roles of the LMs terms are most peripheral.
@@
Figure 8: Differences in the model composition of EMNLP 2020 and 2023, measured respectively on the most LM-focused quarter (${Q4}^{+}$; darker colors) and the least one (${Q1}^{-}$; lighter colors), and ranked by the share of ${Q4}^{+}$ in 23-E. Blue/Red denotes an increase/decrease of share. We annotate beside each bar the exact numbers in 2020 and 2023, separated by ‘/’.
We further measure how the same group changes over time, comparing model compositions at an early conference (EMNLP 2020) versus a most recent one (EMNLP 2023). We depict the change of 15 largest components in Figure 8. The findings are consistent: the most LM-centered group adapts thoroughly to the latest models, while ${Q1}^{-}$ sees a much more modest change, or even increased interest for models pre-2020 like BART Lewis et al. (2020) and COMET Bosselut et al. (2019). Most notable is the contrast for BERT: For ${Q4}^{+}$, the once-dominant component loses as much as 36.8% of $N$, going from 56.0% to 19.2%. However, in ${Q1}^{-}$, the BERT component remains unaffected: it still firmly takes up more than 40% of $N$, and even saw a $\sim$3increase, despite all the new model components and families. This demonstrates that the trend in less LM-centered groups is not merely a moderated or delayed version of that in the top ones, but indeed represents a distinct interpretation and resolution of LMs.
The findings converge to an interesting division: the newest components or those with the latest major models (GPT-3, ChatGPT, LlaMA, and Flan-T5, inter alia; all of which are post-2021) instantly enter the most LM-centered discourse, while the least LM-centered ones persist to favor certain earlier models for a longer period. Thus, we should be aware of the chasm between the co-occurring yet distinct contexts that eventually map to different constructions of the same Ship. For instance, a novel property found with 2023 models as default could be problematic if communicated to a BERT-centered subfield without regard, and further used to justify the use of LMs in a new stake despite wildly different ship drafts and capabilities. In return, the context difference may further hinder communications between groups representing the varied use and interpretations of LMs.